OpenStudy (anonymous):

The area of a trapezoid is 48 square inches. If the bases of the trapezoid are 9 inches and 15 inches, what is the height?

OpenStudy (lgbasallote):

area of trapezoid is \[Area = \frac{h(b_1 + b_2)}{2}\] it is given that area is 48 and bases are 9 and 15 so you have \[48 = \frac{h(9+15)}{2}\] \[48 = \frac{h(24)}{2}\] \[48 = h(12)\] can you solve for h now?

OpenStudy (anonymous):

divide 48 into 12 ?

OpenStudy (lgbasallote):

yes!!! nice